OpenStudy (anonymous):

A shirt company has 3 designs each of which can be made with short or long sleeves. There are 7 color patterns available. How many different types of shirts are available from this company? a. 42 c. 21 b. 12 d. 10

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{3 designs , 2 sleeves , 7 colors}\\~\\ \text{ Types of different shirst }\\~\\ =3\times 2\times 7\\~\\ =42}\end{align}\)

OpenStudy (anonymous):

Wanna help me with more lol

OpenStudy (mathmath333):

yea bring it on

OpenStudy (anonymous):

8C3 * 5C2 Compute the following

OpenStudy (anonymous):

I got like 2 minutes left

OpenStudy (mathmath333):

u want ans or explanation

OpenStudy (anonymous):

I got more questions instead lol

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dbinom{8}{3}\times \dbinom{5}{2}\\~\\~\\ =\dfrac{8!}{3!}\times \dfrac{5!}{2!}\\~\\~\\ =560}\end{align}\)

OpenStudy (mathmath333):

i hope u know how to compute the factorial

OpenStudy (anonymous):

Can I throw some more questions at you? And yea Im kinda getting the grasp of it. Im not sure on it and dont understnad it.

OpenStudy (mathmath333):

what u dont undertand u can ask instantly

OpenStudy (anonymous):

Alot lol.

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{example\\~\\~~8! =8*7*6*5*4*3*2*1\\ }\end{align}\)

OpenStudy (mathmath333):

u can ask a ton

OpenStudy (anonymous):

It wud take a day I swear lol

OpenStudy (mathmath333):

ok then

OpenStudy (anonymous):

If 6 newborn babies are randomly selected, how many different gender sequences are possible? a. 720 c. 12 b. 36 d. 64